OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cosine (5pi/12)

OpenStudy (anonymous):

\[\frac{5\pi}{12}\] is half of \[\frac{5\pi}{6}\]

OpenStudy (anonymous):

then, 5pi/6 = -√3/2 ?

OpenStudy (freckles):

\[\cos^2(\frac{1}{2} \theta)=\frac{1}{2}(1+\cos(\theta)) \\ \text{ you have determined } \theta=\frac{5\pi}{6} \\ \text{ and I think you are actually saying } \cos(\frac{5\pi}{6})=\frac{-\sqrt{3}}{2} \\ \text{ because } \frac{5\pi}{6} \neq \frac{-\sqrt{3}}{2}\]

OpenStudy (anonymous):

would the answer maybe be cos(5pi/12) = √2 - 2√3 /(fraction) 2 ?

OpenStudy (freckles):

what does that say

OpenStudy (anonymous):